OpenStudy (anonymous):

in session 31 : related rates the worked example how he found that r = (3v/5pi)^(1/3) and why when i tried to solve the example with can design method h was different !!?

OpenStudy (phi):

can you describe the problem ? are you asking about the related rates of a sphere with dv/dt= 1000 cc/sec and r= 8 ?

OpenStudy (anonymous):

no it's how we build grain silo

OpenStudy (phi):

ok, I found it. The argument is a bit involved. We fix the volume V (i.e. treat is as a constant). We fix the "shape" cylinder with hemisphere top, and wish to minimize the surface area as a function of two variables: radius of the cylinder (and hemisphere), and h the height of the cylinder. I would start by writing down an expression for the surface area (including the floor): \[ A = 3 \pi r^2 + 2 \pi r h\] we could try to eliminate one of the variables r or h using the "constraint equation" (i.e. expression for the volume). Or we can find the derivative of the Area with respect to one of the variables. Pick r: \[ \frac{dA}{dr} = 6 \pi r \frac{dr}{dr} + 2 \pi \left( r \frac{dh}{dr}+h \frac{dr}{dr} \right)\\ \frac{dA}{dr} = 6 \pi r + 2 \pi \left( r \frac{dh}{dr}+h \right) \] notice we have to treat h as a function of r because r and h are related by a fixed volume: if you change r , for example, and we keep the volume fixed, then the height must change to keep the volume constant.

OpenStudy (phi):

do the "usual thing", set dA/dr = 0 \[ \frac{dA}{dr} = 6 \pi r + 2 \pi \left( r \frac{dh}{dr}+h \right) = 0 \\ 3r +r \frac{dh}{dr}+h = 0 \] now we can solve for r in terms of h if we can get rid of dh/dr

OpenStudy (phi):

to find dh/dr, we use the constraint equation \[ V_0 = \pi r^2 h + \frac{2}{3} \pi r^3 \\ \frac{dV_0}{dr} = 0 = 2 \pi r h + \pi r^2 \frac{dh}{dr} + 2 \pi r^2 \] notice because V0 is a constant, its derivative is 0 solve for dh/dr \[ 2 \pi r h + \pi r^2 \frac{dh}{dr} + 2 \pi r^2 = 0 \\ 2h + r \frac{dh}{dr} +2r = 0\] continuing \[ \frac{dh}{dr} = -2 \frac{r+h}{r} \]

OpenStudy (phi):

use that expression for dh/dr in the expression where we are minimizing the surface area \[ 3r +r \frac{dh}{dr}+h = 0 \\ 3r +-2 \cancel{r} \frac{r+h}{\cancel{r} }+h = 0 \] from which we get \[ 3r -2r -2h+h= 0 \\ r-h= 0 \\ r= h\] this says to minimize the surface area, we want the radius equal the height. use this in the volume expression. i.e. replace h with r and solve for r

OpenStudy (phi):

start with \[ \pi r^2 h + \frac{2}{3} \pi r^3 = V_0 \] and with h= r (to minimize the surface area, we have \[ \frac{3}{3}\pi r^3 + \frac{2}{3} \pi r^3 = V_0 \\ \frac{5}{3} \pi r^3 = V_0 \\ r^3 = \frac{3V_0}{5\pi} \\ r= \left( \frac{3V_0}{5\pi} \right)^\frac{1}{3} \]

OpenStudy (phi):

because V0 is a constant, r has a specific value, and using that value (and the fact that h=r) we can find the minimum value for the surface area: \[ A = 3 \pi r^2 + 2 \pi r h \\ A= 5 \pi r^2 \\ = 5 \pi \left( \frac{3V_0}{5\pi} \right)^\frac{2}{3} \] which can be written in the form \[ A = \sqrt[3]{45 \pi V_0^2} \]